# Joint Graph Rewiring and Feature Denoising via Spectral Resonance

### v2

@@ A.3 Proof of Proposition 1
Notation. We order eigenvalues from largest to smallest and denote the eigenvector associated with the eigenvalue $\lambda_{j}$ by $\bm{v}_{j}\left(\bm{A}\right)$. For the leading eigenvectors of $\bm{A}$ and $\bm{B}$ we write $\bm{v}_{1}\left(\bm{A}\right)=\bm{v}_{A}$ and $\bm{v}_{1}\left(\bm{B}\right)=\bm{v}_{B}$.
@@
Recall that in the BBP transition (Baik et al., 2005), with the random matrices defined in the main text
$\bm{\bm{A}}$ $=\lambda uu^{T}+\bm{O}_{A}$
$\bm{\bm{B}}$ $=\lambda uu^{T}+\bm{O}_{B}$
when $\lambda>1$, we have as $N\to\infty$ that
$\lambda(A)$ $\overset{a.s.}{\rightarrow}\lambda+\frac{1}{\lambda}$ (5)
@@
$\left\langle v_{1}\left(\bm{A}\right),u\right\rangle^{2}$ $\overset{a.s.}{\rightarrow}1-\frac{1}{\lambda^{2}}.$
To prove the proposition, we need to show that when $u$ is replaced by a vector correlated with $\bm{O}_{A}$, the alignment between the spike and the leading eigenvector increases. Based on (5), and if we choose the same direction of the leading eigenvector as of $u$, i.e. $\left\langle u,v_{A}\right\rangle,\left\langle u,v_{B}\right\rangle>0$ (which we can do due to the sign ambiguity in the spectral decomposition), we can write the leading eigenvector of $\bm{A}$ and $\bm{B}$ as
$v_{A}$ $=\sqrt{1-\frac{1}{\lambda^{2}}}u+\frac{1}{\lambda}w_{A}$ (6)
$v_{B}$ $=\sqrt{1-\frac{1}{\lambda^{2}}}u+\frac{1}{\lambda}w_{B},$
where $w_{A},w_{B}$ are two norm-1 vectors orthogonal to $u$, and when $N\to\infty$ we also have $\langle w_{A},w_{B}\rangle\overset{a.s.}{\to}0$. In this case, we can represent the denoised matrix as
$\bm{A}_{\eta}=$ $\lambda uu^{T}+\lambda_{1}(\bm{A})\left(-v_{A}v_{A}^{T}+\tilde{v}_{A}\tilde{v}_{A}^{T}\right)+\bm{O}_{A}$ (7)
$=$ $\lambda uu^{T}+\lambda_{1}(\bm{A})\left(-v_{A}v_{A}^{T}+\left(\left(1-\eta\right)v_{A}+\eta v_{B}\right)\left(\left(1-\eta\right)v_{A}+\eta v_{B}\right)^{T}\right)+\bm{O}_{A}$
$=$ $\lambda_{1}(\bm{S})v_{1}(\bm{S})v_{1}(\bm{S})^{T}+\lambda_{2}(\bm{S})v_{2}(\bm{S})v_{2}(\bm{S})^{T}+\lambda_{3}(\bm{S})v_{3}(\bm{S})v_{3}(\bm{S})^{T}+\bm{O}_{A}$
where $\bm{S}=\lambda uu^{T}+\lambda_{1}(\bm{A})\left(-v_{A}v_{A}^{T}+\left(\left(1-\eta\right)v_{A}+\eta v_{B}\right)\left(\left(1-\eta\right)v_{A}+\eta v_{B}\right)^{T}\right)=\sum_{i=1}^{3}\lambda_{i}(\bm{B})v_{i}(\bm{S})v_{i}(\bm{S})^{T}$ is a rank-3 symmetric matrix, with eigenvalues as functions of $\lambda$ and $\eta$. The eigenvectors are linear combinations of $u$, $w_{A}$ and $w_{B}$ whose weights are also functions of $\lambda$ and $\eta$. In particular, by the eigendecomposition of $\bm{S}$, we get
$\lambda_{1}(\bm{S})$ $=\lambda+c_{\lambda_{1}}\eta^{2}+\mathcal{O}\left(\eta^{3}\right)$ (8)
$\lambda_{2}(\bm{S})$ $=c_{\lambda_{2}}\eta+\mathcal{O}\left(\eta^{2}\right)$
$\lambda_{3}(\bm{S})$ $=c_{\lambda_{3}}\eta+\mathcal{O}\left(\eta^{2}\right)$
$v_{1}(\bm{S})$ $=u+\eta(a_{1}w_{a}+b_{1}w_{b}+c_{1}u)+\mathcal{O}(\eta^{2})=u+\eta\delta u+\mathcal{O}(\eta^{2})$
$v_{2}(\bm{S})$ $=a_{2}w_{a}+b_{2}w_{b}+c_{2}u+\mathcal{O}(\eta)$
$v_{3}(\bm{S})$ $=a_{3}w_{a}+b_{3}w_{b}+c_{3}u+\mathcal{O}(\eta)$
where $c_{\lambda_{1}}$, $c_{\lambda_{2}}$, $c_{\lambda_{3}}$ and $\{a,b,c\}_{\{1,2,3\}}$ are constants that depend only on $\lambda$. In this case, with a small $\eta$, only the leading eigenvalue $\bar{\lambda}:=\lambda_{1}(\bm{S})$ would be outside of the bulk and the corresponding eigenvector is the leading eigenvalue we are looking for. We also note that $c_{\lambda_{1}}>0$, and $c_{\lambda_{2}}+c_{\lambda_{3}}<0$, $a_{2}=-a_{3}$, $b_{1}+a_{1}=0$ and $c_{1}=c_{2}=c_{3}=0$. Using the first-order eigenvector perturbation method compared to the spike-1 matrix $\bar{\bm{A}}=\bar{\lambda}uu^{T}+\bm{O}_{A}$, we have
$\delta v_{1}(\bm{A}_{\eta})$ $:=v_{1}({\bm{A}}_{\eta})-v_{1}(\bar{\bm{A}})$
$=\eta\sum_{j\neq 1}\frac{\lambda_{1}(\bm{S})v_{j}(\bar{\bm{A}})^{T}\left(u\delta u^{T}+\delta uu^{T}\right)v_{1}(\bar{\bm{A}})}{\lambda_{1}\left(\bar{\bm{A}}\right)-\lambda_{j}\left(\bar{\bm{A}}\right)}v_{j}(\bar{\bm{A}})$
$+\eta\sum_{j\neq 1}\frac{v_{j}(\bar{\bm{A}})^{T}\left(c_{\lambda_{2}}v_{2}(\bm{S})v_{2}(\bm{S})^{T}+c_{\lambda_{3}}v_{3}(\bm{S})v_{3}(\bm{S})^{T}\right)v_{1}(\bar{A})}{\lambda_{1}\left(\bar{\bm{A}}\right)-\lambda_{j}\left(\bar{\bm{A}}\right)}v_{j}(\bar{\bm{A}})+\mathcal{O}\left(\eta^{2}\right).$
Recalling (6) and (8), we get
$\left\langle\delta v_{1}({\bm{A}_{\eta}}),u\right\rangle$ $=\lambda\eta\left(\lambda a_{1}+2\left(\lambda-\frac{\lambda}{\lambda^{2}}\right)b_{1}\right)\sum_{j\neq 1}\frac{\left(u^{T}v_{j}\right)^{2}}{\lambda_{1}\left(\bar{\bm{A}}\right)-\lambda_{j}\left(\bar{\bm{A}}\right)}$
$-\eta c_{\lambda_{2}}\sqrt{1-\frac{1}{\lambda^{2}}}a_{2}^{2}\sum_{j\neq 1}\frac{\left(u^{T}v_{j}\right)^{2}}{\lambda_{1}\left(\bar{\bm{A}}\right)-\lambda_{j}\left(\bar{\bm{A}}\right)}$
$-\eta c_{\lambda_{3}}\sqrt{1-\frac{1}{\lambda^{2}}}a_{3}^{2}\sum_{j\neq 1}\frac{\left(u^{T}v_{j}\right)^{2}}{\lambda_{1}\left(\bar{\bm{A}}\right)-\lambda_{j}\left(\bar{\bm{A}}\right)}+\mathcal{O}\left(\eta^{2}\right)$
$=\eta\left(-\lambda a_{1}\left(\lambda-\frac{2}{\lambda}\right)-(c_{\lambda_{2}}+c_{\lambda_{3}})a_{2}^{2}\sqrt{1-\frac{1}{\lambda^{2}}}\right)\sum_{j\neq 1}\frac{\left(u^{T}v_{j}\right)^{2}}{\lambda_{1}\left(\bar{\bm{A}}\right)-\lambda_{j}\left(\bar{\bm{A}}\right)}+\mathcal{O}\left(\eta^{2}\right)$
Noting that $a_{1}<0$ and $c_{\lambda_{2}}+c_{\lambda_{3}}<0$, for any $\lambda>\sqrt{2}$, we can always find a $\eta=\eta^{0}$, which makes $\langle\delta v_{1}({\bm{A}}_{\eta}),u\rangle$ positive. Since $\bar{\bm{A}}$ follows the standard BBP transition, we finally have
$\left\langle v_{1}({\bm{A}}_{\eta}),u\right\rangle=\left\langle v_{1}(\bar{\bm{A}}),u\right\rangle+\left\langle\delta v_{1}({\bm{A}_{\eta}}),u\right\rangle>\sqrt{1-\frac{1}{\bar{\lambda}^{2}}}>\sqrt{1-\frac{1}{{\lambda}^{2}}}.$
The above results are primarily based on the eigendecomposition of $\bm{B}$, for which analytical expressions can be derived, but they are quite complex. We can easily verify the results for specific values of $\lambda$. For $\lambda=2$ we have
$c_{\lambda_{1}}=\frac{75}{64},c_{2}=-\frac{5\left(\sqrt{2}+1\right)}{8},c_{3}=\frac{5\left(\sqrt{2}-1\right)}{8},a_{1}=-\frac{5\sqrt{3}}{16},a_{2}=\frac{8\sqrt{\frac{2}{3}}}{5}$
and for $\lambda=4$ we have
$c_{\lambda_{1}}=\frac{4335}{8192},c_{2}=-\frac{17\left(\sqrt{2}+1\right)}{64},c_{3}=\frac{17\left(\sqrt{2}-1\right)}{64},a_{1}=-\frac{17\sqrt{15}}{256},a_{2}=\frac{128\sqrt{\frac{2}{15}}}{17}.$
@@ A.4 Datasets
Table 5 shows the properties of the real-world datasets used. We also provide the homophily measure $\mathcal{H}(\mathcal{G})$ proposed in Pei et al. (2020), which we compute using the build-in function of Pytorch Geometric (Fey and Lenssen, 2019). For the cSBM, following (Chien et al., 2021), we choose $N=5000$, $F=2000$ and thus have $\gamma=\frac{N}{F}=2.5$. Since the threshold to recover communities in cSBM is $\lambda^{2}+\mu^{2}/\gamma>1$ (Deshpande et al., 2018), we use a margin such that $\lambda^{2}+\mu^{2}/\gamma=1+\epsilon$. We choose the same $\epsilon=3.25$ as Chien et al. (2021) in all our experiments to be above the detection threshold and $d=5$ to obtain a sparse graph to be close to the properties of real-world graphs. From the recovery threshold, we can parameterize the resulting arc of an ellipse with $\lambda\geq 0$ and $\mu\geq 0$ using $\phi=\arctan(\lambda\sqrt{\gamma}/\mu)$. Table 6 shows the parameters $\mu^{2}$ and $\lambda$ and the homophily measure $\mathcal{H}(\mathcal{G})$for the different values of $\phi$.
@@ A.5 Additional Results
We provide a number of additional experiments which did not fit in the main text. These include more experiments on additional heterophilic datasets from Platonov et al. (2023), results for the homophilic datasets in the dense splitting, more experiments with DIGL (Gasteiger et al., 2019), more alignment results and results for synthetic and real-world data using spectral clustering with and w/o JDR. The clustering experiments in particular allow an interpretation of how JDR works: Applying it to a graph increases its “spectral clusterability”.
@@ A.5.1 Additional Heterophilic Datasets
In order to get a more comprehensive picture of the performance of JDR, we also test JDR on all the datasets proposed there by Platonov et al. (2023). Table 7 shows the results on these datasets using their original splits and comparing DIGL, FoSR, BORF and JDR. In general, the performance increases are relatively small for all methods, but overall JDR still performs best. It achieves significant performance increases on Tolokers and Questions. For Minesweeper none of the methods is really able to improve performance. The reason for this is in the synthetic design of its graph and the features: The graph does not contain any information about the labels, as it only connects neighboring cells (it is solely a computational graph). The same is partially true for node features which indeed contain information about neighboring mines, but only for $50\%$ of the nodes. This renders JDR unsuitable, which is also reflected in an interesting way in the experiments: We found that the choice of hyperparameters has hardly any influence on performance. But also the results of the other rewiring methods indicate that they cannot be applied here. The graph is a standard grid-graph, which should not exhibit any interesting geometric properties and not contain any insights about the labels. Overall, this is a typical error case for any rewiring method. But it could also be discussed to what extent this dataset is an interesting graph dataset for node classification at all, since the connectivity does not contain any information about the labels.
Table 7: Comparison of DIGL, FoSR, BORF and JDR on real-world heterophilic datasets from Platonov et al. (2023): Mean accuracy (%) and ROC AUC for imbalanced Minesweeper, Tolokers and Questions± $\pm\ 95\%$ confidence interval. Best average accuracy in bold. OOM indicates an out-of-memory error.
@@ A.5.2 Homophilic Datasets in the Dense Splitting
A.5.3 Combining JDR and DIGL

### v3

@@ A.3 Proof of Proposition 1
Notation. We order the eigenvalues and singular values from largest to smallest and denote the eigenvector associated with the eigenvalue $\lambda_{j}$ by $\bm{v}_{j}\left(\bm{A}^{c}\right)$. For the leading eigenvalue and eigenvectors of $\bm{A}^{c}$, we write $\lambda_{1}=\lambda_{A}$ and $\bm{v}_{1}\left(\bm{A}^{c}\right)=\bm{v}_{A}$. We use analogous notation for the singular values and corresponding singular vectors of $\bm{X}$. For simplicity and without loss of generality we assume that the angles between these vectors are accute, i.e., $\left\langle\bm{v}_{A},\bm{u}_{X}\right\rangle,\left\langle\tilde{\bm{y}},\bm{v}_{A}\right\rangle,\left\langle\tilde{\bm{y}},\bm{u}_{X}\right\rangle\geq 0$.
@@
When $\lambda>1$, based on the Baik-Ben Arous-Péché (BBP) transition [Baik et al., 2005, Paul, 2007], the leading eigenvalue of $\bm{A}$ lies outside the spectral bulk,
$\lambda_{A}=\lambda+\frac{1}{\lambda}+\mathcal{O}_{p}\left(\frac{1}{\sqrt{N}}\right),$
and the fluctuation of the leading eigenvector satisfies
[…]
where $\mathrm{Haar}(\mathbb{S}_{\tilde{\bm{y}}^{\perp}}^{N-2})$ is the uniform distribution on the sphere orthogonal to $\tilde{\bm{y}}$, $S_{\tilde{\bm{y}}^{\perp}}^{n-2}=\left\{\bm{v}:\bm{v}\in\mathrm{\mathbb{R}}^{N}\mid\bm{v}^{T}\tilde{\bm{y}}=0,\left\|\bm{v}\right\|=1\right\}$, and the convergence is in distribution as $N\to\infty$. Similarly, for the rectangular matrix $\bm{X}$ we have [Benaych-Georges and Nadakuditi, 2012],
$\sigma_{X}=\sqrt{\frac{\left(\gamma+\mu\right)\left(1+\mu\right)}{\mu}}+\mathcal{O}_{p}\left(\frac{1}{\sqrt{N}}\right)$
and the fluctuation of the leading singular vector satisfies
[…]
To denoise $\bm{A}^{c}$, we adjust the leading eigenvector towards the direction of $\bm{u}_{X}$ as
$\tilde{\bm{v}}_{A}=\left(1-\eta_{A}\right)\bm{v}_{A}+\eta_{A}\bm{u}_{X}$
where $\eta_{A}>0$ is a small constant. The corresponding perturbation in the matrix reads
$\bm{A}^{c}_{\eta_{A}}-\bm{A}^{c}$ $=\lambda_{A}(\tilde{\bm{v}}_{A}{\tilde{\bm{v}}_{A}}^{T}-\bm{v}_{A}\bm{v}_{A}^{T})$
$=\lambda_{A}\eta_{A}\left(\bm{v}_{A}\bm{u}_{X}^{T}+\bm{u}_{X}\bm{v}_{A}^{T}\right)+\mathcal{O}\left(\eta_{A}^{2}\right)$
$=\lambda_{A}\eta_{A}\bm{E}+\mathcal{O}\left(\eta_{A}^{2}\right).$
The first-order perturbation of the leading eigenvector yields
$\bm{v}_{1}{(\bm{A}^{c}_{\eta_{A}})}-\bm{v}_{1}{(\bm{A}^{c})}=\lambda_{A}\eta_{A}\sum_{j\neq 1}\frac{\bm{v}_{j}(\bm{A}^{c})^{T}\left(\bm{v}_{A}\bm{u}_{X}^{T}+\bm{u}_{X}\bm{v}_{A}^{T}\right)\bm{v}_{A}}{\lambda_{A}-\lambda_{j}}\bm{v}_{j}(\bm{A}^{c})+\mathcal{O}\left(\eta_{A}^{2}\right).$
Since $\bm{v}_{j}(\bm{A}^{c})^{T}\bm{v}_{A}=0$ for $j>1$, we have
$\left\langle\bm{v}_{1}{(\bm{A}^{c}_{\eta_{A}})}-\bm{v}_{1}{(\bm{A}^{c})},\tilde{\bm{y}}\right\rangle$
$=\lambda_{A}\eta_{A}\sum_{j\neq 1}\frac{\tilde{\bm{y}}^{T}\bm{v}_{j}(\bm{A}^{c})\bm{v}_{j}(\bm{A}^{c})^{T}\bm{u}_{X}\bm{v}_{A}^{T}\bm{v}_{A}}{\lambda_{A}-\lambda_{j}}+O\left(\eta_{A}^{2}\right)$
$=\lambda_{A}\eta_{A}c_{1}\sum_{j\neq 1}\frac{\left(\tilde{\bm{y}}^{T}\bm{v}_{j}(\bm{A}^{c})\right)^{2}}{\lambda_{A}-\lambda_{j}}+\lambda_{A}\eta_{A}c_{2}\sum_{j\neq 1}\frac{\tilde{\bm{y}}^{T}\bm{v}_{j}(\bm{A}^{c})\bm{v}_{j}(\bm{A}^{c})^{T}\bm{q}_{X}}{\lambda_{A}-\lambda_{j}}+\mathcal{O}\left(\eta_{A}^{2}\right)$ (6)
[…]
$\frac{1}{\frac{1}{\lambda}+\lambda+2}+\mathcal{O}_{p}\left(N^{-\frac{2}{3}}\right)<\frac{1}{\lambda_{A}-\lambda_{j}}<\frac{1}{\frac{1}{\lambda}+\lambda-2}+\mathcal{O}_{p}\left(N^{-\frac{2}{3}}\right)\quad\text{for}\quad j>1.$
Therefore the first term in (6) can be bounded as $\sum_{j\neq 1}\frac{\left(\tilde{\bm{y}}^{T}\bm{v}_{j}(\bm{A}^{c})\right)^{2}}{\lambda_{A}-\lambda_{j}}\overset{a.s}{>}\frac{1}{\lambda_{A}^{2}}\frac{1}{\frac{1}{\lambda}+\lambda+2}$ when $N\to\infty$. For the second term, we note that the vector $\bm{q}_{X}$ is independent of $\bm{z}:=\sum_{i\neq 1}\frac{\tilde{\bm{y}}^{T}\bm{v}_{j}(\bm{A}^{c})}{\lambda_{A}-\lambda_{j}}\bm{v}_{j}(\bm{A}^{c})$. Each eigenvalue $\bm{v}_{j}(\bm{A}^{c})$ is uniformly distributed on $\mathbb{S}_{\bm{v}_{A}^{\perp}}^{N-2}$ and $\left\{\bm{v}_{j}(\bm{A}^{c})\right\}_{j=2}^{N}$ is an orthogonal basis of $\mathbb{R}_{\bm{v}_{A}^{\perp}}^{N-1}=\{\bm{v}:\bm{v}\in\mathbb{R}^{N}\mid\bm{v}^{T}\bm{v}_{A}=0\}$. Therefore, for large $N$, $\tilde{\bm{y}}^{T}\bm{v}_{j}(\bm{A}^{c})$ is approximately independent of each element in $\bm{v}_{j}(\bm{A}^{c})$. More precisely, the entries of $\bm{z}$ are of the order of $\mathcal{O}_{p}\left(\frac{1}{\sqrt{N}}\right)$, and thus $\sum_{j\neq 1}\frac{\tilde{\bm{y}}^{T}\bm{v}_{j}(\bm{A}^{c})\bm{v}_{j}(\bm{A}^{c})^{T}\bm{q}_{X}}{\lambda_{A}-\lambda_{j}}=\left\langle\bm{q}_{X},\bm{z}\right\rangle=\mathcal{O}_{p}\left(\frac{1}{\sqrt{N}}\right)$. Summarizing, we get
@@
$\left\langle\bm{v}_{1}{(\bm{A}^{c}_{\eta_{A}})}-\bm{v}_{1}{(\bm{A}^{c})},\tilde{\bm{y}}\right\rangle\overset{a.s.}{>}\frac{\sqrt{1-\frac{\gamma\left(1+\mu\right)}{\mu\left(\mu+\gamma\right)}}}{\left(\lambda+1/\lambda\right)\left(\lambda+1/\lambda+2\right)}\eta_{A}+\mathcal{O}\left(\eta_{A}^{2}\right)\quad\text{when}\quad N\to\infty.$
A similar strategy can be used to show that $\left\langle\bm{u}_{1}({\bm{X}_{\eta_{X}}}),\tilde{\bm{y}}\right\rangle\overset{a.s}{>}\left\langle\bm{u}_{1}({\bm{X}}),\tilde{\bm{y}}\right\rangle$. ∎
@@ A.4 Datasets
Table 6 shows the properties of the real-world datasets used. We also provide the homophily measure $\mathcal{H}(\mathcal{G})$ proposed in Pei et al. [2020], which we compute using the build-in function of Pytorch Geometric [Fey and Lenssen, 2019]. For the cSBM, following [Chien et al., 2021], we choose $N=5000$, $F=2000$ and thus have $\gamma=\frac{N}{F}=2.5$. Since the threshold to recover communities in cSBM is $\lambda^{2}+\mu^{2}/\gamma>1$ [Deshpande et al., 2018], we use a margin such that $\lambda^{2}+\mu^{2}/\gamma=1+\epsilon$. We choose the same $\epsilon=3.25$ as Chien et al. [2021] in all our experiments to be above the detection threshold and $d=5$ to obtain a sparse graph to be close to the properties of real-world graphs. From the recovery threshold, we can parameterize the resulting arc of an ellipse with $\lambda\geq 0$ and $\mu\geq 0$ using $\phi=\arctan(\lambda\sqrt{\gamma}/\mu)$. Table 7 shows the parameters $\mu^{2}$ and $\lambda$ and the homophily measure $\mathcal{H}(\mathcal{G})$for the different values of $\phi$.
@@ A.5 Additional Results
We provide a number of additional experiments which did not fit in the main text. These include more experiments on additional heterophilic datasets from Platonov et al. [2023], results for the homophilic datasets in the dense splitting, more experiments with DIGL [Gasteiger et al., 2019], more alignment results and results for synthetic and real-world data using spectral clustering with and w/o JDR. The clustering experiments in particular allow an interpretation of how JDR works: Applying it to a graph increases its “spectral clusterability”.
@@ A.5.1 Additional Heterophilic Datasets
In order to get a more comprehensive picture of the performance of JDR, we also test JDR on all the datasets proposed there by Platonov et al. [2023]. Table 8 shows the results on these datasets using their original splits and comparing DIGL, FoSR, BORF and JDR. In general, the performance increases are relatively small for all methods, but overall JDR still performs best. It achieves significant performance increases on Tolokers and Questions. For Minesweeper none of the methods is really able to improve performance. The reason for this is in the synthetic design of its graph and the features: The graph does not contain any information about the labels, as it only connects neighboring cells (it is solely a computational graph). The same is partially true for node features which indeed contain information about neighboring mines, but only for $50\%$ of the nodes. This renders JDR unsuitable, which is also reflected in an interesting way in the experiments: We found that the choice of hyperparameters has hardly any influence on performance. But also the results of the other rewiring methods indicate that they cannot be applied here. The graph is a standard grid-graph, which should not exhibit any interesting geometric properties and not contain any insights about the labels. Overall, this is a typical error case for any rewiring method. But it could also be discussed to what extent this dataset is an interesting graph dataset for node classification at all, since the connectivity does not contain any information about the labels.
Table 8: Comparison of DIGL, FoSR, BORF and JDR on real-world heterophilic datasets from Platonov et al. [2023]: Mean accuracy (%) and ROC AUC for imbalanced Minesweeper, Tolokers and Questions± $\pm\ 95\%$ confidence interval. Best average accuracy in bold. OOM indicates an out-of-memory error.
@@ A.5.2 Homophilic Datasets in the Dense Splitting
A.5.3 MLP with JDR
We design JDR with the aim to denoise the (possibly) complementary information in graph and features. This is based on the claim that a GNN is the method of choice when both the graph and the features contain valuable information, as it can utilize both. The experiments on cSBM in the main text show the ability of JDR to transfer information between the two. This becomes visible especially in the corner cases, were either the graph or the features do not contain any information about the labels. In this case the only way to improve is by transferring label information from one source to the other. To investigate this further, we test an MLP with JDR and compare the results with only the MLP and the GNNs. Since JDR can transfer the information from the graph to the features, an MLP should be able to perform similar to the GNNs. Therefore, we tune the hyperparameters of JDR with the MLP downstream model on the synthetic cSBM datasets and the real-world datasets from the main text.
The results on the cSBM data are shown in Figure 6. They show that combining an MLP with JDR clearly outperforms GCN in the heterophilic regime and performs very similar to GPRGNN. If the GNNs are also combined with JDR, they generally again provide superior performance compared to MLP+JDR. In the weak graph regime they have a huge performance advantage of about $20\%$, since JDR cannot improve the feature quality for the MLP in these cases. In the heterophilic regime, MLP+JDR is still comparable to GCN+JDR, which is not the case for the homophilic regime. A similar behavior (without JDR) has been observed in the literature before, e.g. by Ma et al. [2022] and is related to the limited ability of GCN to deal with heterophilic graphs. GPRGNN, however, does not show this limitation and provides superior performance across all datasets. Notably, all these findings do directly translate to the real-world datasets. The results on the real-world data can be found in Table 10 in the homophilic case and in Table 11 for the heterophilic case. For the homophilic datasets, MLP+JDR shows clear performance increases but cannot beat any of the GNN or GNN+JDR baselines. On the heterophilic graph, the MLP already outperforms GCN on three out of five datasets and with JDR on all datasets. GCN+JDR regains superior performance on three of the five datasets, but only GPRGNN+JDR outperforms the MLP+JDR on all datasets. Overall, this supports the claim that a GNN like GPRGNN is the method of choice when both the graph and the features contain valuable information.
Figure 6: Comparison of MLP and the GNNs on the cSBM datasets in the sparse splitting. Comparisons of MLP and MLP+JDR with the GNNs (7(a)) and GNN+JDR (7(b)). The error bars indicate the $95\%$ confidence interval. The MLP outperforms the GNNs in the very weak graph regime. Combining it with JDR clearly beats GCN (especially in the heterophilic regime) and performs very similar to GPRGNN. If the GNNs are combined with JDR, they generally again provide superior performance compared to MLP+JDR.
Table 10: Results of MLP and JDR on real-world homophilic dataset using the sparse splitting: Mean accuracy (%) ± $\pm\ 95\%$ confidence interval. Best average accuracy in bold.
Method Cora CiteSeer PubMed Computers Photo
MLP $50.79$±$\pm 0.73$ $50.29$±$\pm 0.48$ $79.73$±$\pm 0.13$ $73.17$±$\pm 0.31$ $80.88$±$\pm 0.33$
MLP+JDR $62.66$±$\pm 0.61$ $61.55$±$\pm 0.32$ $80.86$±$\pm 0.12$ $80.65$±$\pm 0.24$ $88.34$±$\pm 0.45$
Table 11: Results of MLP and JDR on real-world heterophilic dataset using the dense splitting: Mean accuracy (%) ± $\pm\ 95\%$ confidence interval. Best average accuracy in bold.
Method Chameleon Squirrel Actor Texas Cornell
MLP $49.07$±$\pm 0.57$ $28.19$±$\pm 0.40$ $38.54$±$\pm 0.30$ $91.16$±$\pm 0.79$ $88.19$±$\pm 0.74$
MLP+JDR $70.48$±$\pm 0.46$ $59.18$±$\pm 0.31$ $39.50$±$\pm 0.26$ $91.16$±$\pm 0.77$ $88.47$±$\pm 0.77$
GCN+JDR $69.76$±$\pm 0.50$ $\bm{61.76}$±$\pm\bm{0.39}$ $40.47$±$\pm 0.31$ $85.12$±$\pm 0.74$ $84.51$±$\pm 1.06$
GPRGNN+JDR $\bm{71.00}$±$\pm\bm{0.50}$ $60.62$±$\pm 0.38$ $\bm{41.89}$±$\pm\bm{0.24}$ $\bm{93.85}$±$\pm\bm{0.54}$ $\bm{89.45}$±$\pm\bm{0.84}$
Figure 7: Comparison of DIGL [Gasteiger et al., 2019] and JDR on the cSBM datasets in the sparse splitting. Results for (7(a)) GCN and (7(b)) GPRGNN as downstream models. The error bars indicate the $95\%$ confidence interval. As expected, DIGL is not really able to improve the performance of the GNNs in the heterophilic regime. It achieves the greatest improvement in the weak-graph regime and for strongly homophilic graphs, especially using GCN as downstream model. Another interesting observation is that for GCN and $\phi<0.25$ the curve of MLP corresponds exactly to the one of GCN +DIGL. The reason for this is that the hyperparameters found for DIGL ensure that the graph is ignored ($\alpha=1.0$), which means that the GCN then collapses to a simple MLP. For the more powerfull GPRGNN, on the other hand, DIGL is generally hardly able to improve performance, while JDR clearly increases the performance across all $\phi$.
A.5.4 Combining JDR and DIGL